# Newcastle-upon-Tyne, UK

### v3

@@ 1 Introduction
Modern mobile devices come equipped with an array of embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This rich sensor data has enabled applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] or reduce user-driven authentication prompts [7, 8, 9, 10, 11]. Despite these advancements, fundamental questions remain. Multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors have not been rigorously evaluated. Many existing studies rely on heuristic assessments or machine learning classifiers, e.g. [4, 13, 14, 2, 12], that do not address core security requirements: how much entropy do sensors truly provide, and whether multi-modal combinations yield proportional security gains or suffer from inter-sensor correlations.
Our analysis reveals systemic limitations. Commodity sensors exhibit significant biases, with gyroscopes yielding min-entropy as low as 0.96 bits—well below cryptographic thresholds. Although multi-modal sensor fusion mitigates some of these issues, non-uniform distributions and inter-sensor correlations reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy. These findings challenge the notion that increasing the number of sensors automatically strengthens security, and underscore the inadequacy of raw sensor data as a reliable entropy source. Our contributions are:
@@
The rest of this paper is organised as follows: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis framework is made publicly available to foster future research. [^1] https://github.com/cgshep/entropy-collapse-mobile-sensors
@@ 2.1 Mobile Sensor-based Security Applications
One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [15, 13, 14, 4]. Mehrnezhad et al.[13] introduced a technique that uses accelerometer readings to verify that an NFC payment instrument and terminal are physically tapped together. The authors posit that “physical tapping causes random but correlated vibrations at both devices, which are hard to forge (or reproduce)” (p.1, [13]). The work reports an equal error rate (EER) of 17.65% using a machine learning-based approach. Gurulian et al. [14] also explored the use of shared vibration patterns between users using unique vibration patterns generated by one device. Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [16, 17, 18, 8, 9, 19, 10, 11], where sensor data is used to authenticate users passively without explicit interaction. (with up to 99% classification accuracy in some instances [17]). Mobile sensors have also been proposed as enabling novel device pairing schemes for secure communication, e.g. [1, 20].
@@
The notion of ‘hardness’ is typically inferred based on a machine learning-based analysis; it is not grounded in standard quantitative methods used in the field of authentication, such as Shannon [21] and min-entropy [22]. A general approach to follows one of those given by Shrestha et al. [3]. Here, sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13]. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] have been used to evaluate the effectiveness of distinguishing between legitimate and illegitimate samples. (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].) We observe that machine learning-based evaluation metrics do not directly measure entropy, which is a crucial factor in determining whether sensor data is a suitable source of randomness for security applications. Some work has attempted to assess the entropy of sensor signals, but this represents a minority. T2Pair by Li et al. [23] 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al. [24] reporting 51–54 bits of Shannon entropy. Even in the best cases, this is low relative to modern cryptographic standards.
We point to adjacent work that has questioned the utility of mobile sensors. Markantonakis et al. [4], building on Gurulian et al. [5] and Shepherd et al. [6], presented a reproducibility study disputing the claimed effectiveness of mobile sensors when deployed under real-world time constraints. Using consumer off-the-shelf devices, the work reports that sensors available through the Android platform could not be used to reliably detect man-in-the-middle attacks in NFC contactless transactions within the 500ms time window required by the EMV payment protocol. Under those constraints, the effectiveness drastically decreased compared to previously reported results. Indeed, it was lower to the extent such that sensors could not be recommended for mobile transactions without posing significant usability and security issues in practice.
@@ 2.2 Sensor Entropy Analyses
In earlier work, Voris et al. [22] investigated the use of accelerometers as true random number generators (TRNGs), specifically aboard a WISP RFID tag and a Nokia N97 mobile phone. The authors find that min-entropy—defined in §2.3—that min-entropy is proportional to the motion applied to the device, and that stationary movement has the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC) are considered other significant influences on the generated entropy. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [25] analyse three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method used. Krhovják et al. [7] examine the entropy of image and audio data collected from mobile phone cameras and microphones respectively. Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropy values of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [26] present an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. The work uses a single-sensor analysis, presenting min-entropy values between 0–7.85. Motion sensors, e.g. accelerometer and vibration sensors, yielded the highest entropy.
@@
Mobile sensors have also been proposed as entropy sources in low-cost RNG designs. Suciu et al. [27] propose using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors. Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [28] suite, but no precise entropy values are presented. Wallace et al. [29] explore the use of accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources. Mobile sensors are inherently opportunistic entropy sources, contrasting with dedicated hardware like ring oscillators, those build on Johnson-Nyquist thermal noise, or quantum-based generators (e.g. see [30]). Mobile sensors depend heavily on user behaviour and environmental, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based security mechanisms largely overlook these dynamics, relying instead on heuristic assessments or those based on machine learning-based evaluation metrics [8, 11]. Such approaches do not account well for skewed distributions and other biases in the underlying data. This contrasts with typical entropy measures used in RNG analyses or in the study of authentication systems [31]. For instance, NIST SP800-90B [32] and AIS 20/31 [33] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@ 2.3 Definitions
We draw attention to four special cases of $\alpha$ that are widely used in the literature. Firstly, the Hartley (Max) Entropy ($\alpha=0$), given in Eq. 2, is the logarithm of the number of possible outcomes that have non-zero probability; it serves effectively as an upper bound.
$H_{0}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
@@
$H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy is thus used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes. This is the Min-Entropy (Eq. 5), i.e. the value at the limit $\alpha\to\infty$.
$H_{\infty}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
@@ 3.1 Threat Model and Assumptions
We consider an adversary whose primary objective is to compromise sensor-based key generation or authentication schemes by exploiting the inherent biases or predictable patterns within sensor data. In particular, we assume the adversary has full or near-full knowledge of the probability distribution of sensor outputs. They may obtain this through statistical modeling of publicly available datasets, prolonged observation of target users, or eavesdropping on sensor data streams.
We exclude adversarial capabilities such as fault-injection attacks or hardware attacks, e.g. see [34]. While such attacks could further reduce the effective entropy space, we regard them as a higher level of sophistication—an important future research direction but not the focus here. Lastly, rather than brute-forcing all possible sensor outputs, we assume a more strategic approach: the adversary uses its knowledge of the sensor’s distribution to prioritise the most probable (or most easily predicted) outcomes. By iterating guesses in order of descending probability, they minimise their expected effort. This threat model represents a realistic adversary who can capitalise on statistical biases in sensor data but does not directly sabotage or physically compromise the device. Our goal is to evaluate whether typical sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist such knowledge-driven attacks.
@@ 3.2 High-level Methodology
    We merge sensor readings into a single, global distribution for each sensor modality in each dataset. For sensors that are inherently discrete or quantised (e.g., integer output ranges), we simply count occurrences. For sensors that produce (quasi-)continuous values, we rely on quantisation strategies (Freedman–Diaconis binning) to partition the output space and approximate an empirical probability mass function.
@@
  • University of Sussex–Huawei Locomotion (SHL) [40, 41]: sampled at 100 Hz. Multiple Huawei Mate 9 phones attached across the users’ body in a body area network (BAN). The publicly available SHL Preview dataset is used, comprising three recording-days per user (a total of 59 hours of data). To scope this study, we use the motion sensors dataset from the handheld mobile phone as a good fit with related work (see §2).
@@ 4 Entropy Analysis
In this section, we analyze the intrinsic entropy of sensor data under the threat model described in §3. We begin by discussing the challenges in quantising naturally continuous sensor values for discrete-entropy calculations, then present our single-sensor findings. Later sections examine multi-sensor joint entropies and the complexities of sensor fusion.
@@ 4.1 Pre-processing
A crucial, yet underexplored, issue in prior work (e.g. [22, 25, 26, 31]) is how to convert inherently continuous sensor outputs into suitable discrete values for entropy estimation. For example, Shannon entropy and min-entropy, as defined in Eqs. 3 and 5, rely on discrete random variables. Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution. However, a sensor’s advertised resolution (e.g. 12 bits for the widely used Bosch BMA mobile accelerometer [43]) does not imply uniform or truly high-fidelity coverage across its range. Everyday usage introduces biases and clustering, resulting in some measurements occurring far more frequently than others. For instance, UCI-HAR data shows accelerometer readings widely spread in certain axes, while approximately 60% of gyroscope readings hover near zero (see Figure 2). Such heavy skew or near-constant values radically diminish real-world entropy, which would otherwise benefit from uniformly distributed values.
@@ 4.2 Single Sensors
Given the biases discussed above, it is unsurprising that some sensor readings exhibit high predictability in the worst-case sense. We illustrate this by calculating individual-sensor entropy across multiple datasets, splitting multi-dimensional readings (e.g., triaxial accelerometer or gyroscope) into separate axes following Voris et al. [22]. The results are given in Table 1. We note that in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors. Rather, the authors preprocess triaxial data into its vector magnitude, i.e. $\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$. We give this as “X.Mag” for a given sensor X. For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this abstract sensor modality.
@@
Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data for that modality and dataset.

### v4

@@ 1 Introduction
Modern mobile devices come equipped with an array of embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This rich sensor data has enabled applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] and reduce user authentication prompts [7, 8, 9, 10, 11]. Despite these advancements, fundamental issues remain: multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated. Many existing studies rely on heuristic assessments or machine learning classifiers, e.g. [4, 13, 14, 2, 12], that do not address critical security questions. That is, firstly, how much entropy do sensors truly provide? And, secondly, to what extent do multi-modal sensor combinations provide security gains?
Our analysis reveals systemic limitations. Commodity sensors exhibit significant biases, with gyroscopes yielding min-entropy as low as 0.96 bits. While multi-modal sensor usage somewhat addresses these issues, non-uniform distributions and inter-sensor correlations reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy. These findings challenge the notion that increasing the number of sensors reliably strengthens security, and it underscores the inadequacy of using raw sensor data as a dependable entropy source. Our contributions are as follows:
@@
The rest of this paper is organised as such: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis work is released publicly to foster future research. [^1] https://github.com/cgshep/entropy-collapse-mobile-sensors
@@ 2.1 Mobile Sensor-based Security Applications
One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [15, 13, 14, 4, 2]. Mehrnezhad et al.[13] introduced a technique that uses accelerometer readings to verify that an NFC payment instrument and terminal are physically tapped together. The authors posit that “physical tapping causes random but correlated vibrations at both devices, which are hard to forge (or reproduce)” (p.1, [13]). The work reports an equal error rate (EER) of 17.65% using a machine learning-based approach. Gurulian et al. [14] also explored the use of shared vibration patterns between users using unique vibration patterns generated by one device. Shrestha et al. [2] explore four environmental modalities, including temperature and humidity, for proximity detection. In this work, single sensors yields 0.733–0.881 F1-score, while combining multiple increases the overall performance to 0.913–0.957. Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [16, 17, 18, 8, 9, 19, 10, 11], where sensor data is used to authenticate users passively without explicit interaction (with up to 99% accuracy claimed in some work [17]). Sensors have also been used to underpin the security of novel device pairing schemes, whether as a primary [20] or second line of authentication [1]. These schemes generally rely on detecting similar motion patterns between two devices using joint accelerometer and gyroscope sensor measurements in order to provide evidence of co-location.
@@
The notion of ‘hardness’ is typically inferred using model evaluation metrics; it is not grounded in standard methods used in the field of authentication, such as Shannon [21] and min-entropy [22]. A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2]. Features are subsequently classified using simple threshold-based or supervised classification models, e.g. Support Vector Machine (SVM) and Random Forest. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are then used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples. (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [18, 1].)
We observe that the vast majority of proposals do not directly measure entropy, which is a crucial factor in determining whether sensor data cannot be easily predicted by an adversary. Some work has attempted to assess the entropy of sensor signals, but this represents a minority. T2Pair by Li et al. [23] yielded 32.3–38.5 bits of Shannon entropy, with a refinement by Wu et al. [24] reporting 51–54 bits of Shannon entropy. Even in the best cases, this is low relative to modern cryptographic standards. [^2] See AIS 20/31 [25]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator. Lastly, we point to adjacent work that has questioned the utility of mobile sensors. Markantonakis et al. [4], building on Gurulian et al. [5] and Shepherd et al. [6], presented a reproducibility study disputing the claimed effectiveness of mobile sensors when deployed under real-world time constraints for NFC-based contactless transactions. Using a 500ms window for data acquisition, as stipulated by the EMV payment protocol, 0.179–0.246 EER was reported depending on the given sensor combination. Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
@@ 2.2 Sensor Entropy Analyses
In earlier work, Voris et al. [22] investigated the use of accelerometers as true random number generators (TRNGs), specifically aboard a WISP RFID tag and a Nokia N97 mobile phone. The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, and that stationary movement has the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC) are considered other significant influences on the generated entropy. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [26] analyse three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method used. Krhovják et al. [7] examine the entropy of image and audio data collected from mobile phone cameras and microphones respectively. Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropy values of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [27] present an analysis of 10 sensor modalities available on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. The work uses a single-sensor analysis, presenting min-entropy values between 0–7.85. Motion sensors, e.g. accelerometer and vibration sensors, yielded the highest entropy.
@@
Mobile sensors have also been proposed as entropy sources in low-cost RNG designs. Suciu et al. [28] propose using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors. Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [29] suite, but no precise entropy values are presented. Wallace et al. [30] explore the use of accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources. Mobile sensors are inherently opportunistic entropy sources, contrasting with dedicated hardware like ring oscillators, those build on Johnson-Nyquist thermal noise, or quantum-based generators (e.g. see [31]). Mobile sensors depend heavily on user behaviour and environmental, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based security mechanisms largely overlook these dynamics, relying instead on heuristic or model evaluation metrics [8, 11]. Such approaches do not account well for skewed distributions and other biases in the underlying data. This contrasts with typical entropy measures used in RNG analyses or in the study of authentication systems [32]. For instance, NIST SP800-90B [33] and AIS 20/31 [25] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@ 2.3 Definitions
We draw attention to four special cases of $\alpha$ that are widely used in the literature. Firstly, the Hartley (Max) Entropy, given in Eq. 2, is the logarithm of the number of possible outcomes that have non-zero probability; it serves effectively as an upper bound.
$H_{0}(X)\equiv\lim_{\alpha\to 0}H_{\alpha}(X)\,=\,\log\Bigl{|}\bigl{\{}\,x\in\mathcal{X}:p(x)>0\bigr{\}}\Bigr{|}.$ (2)
@@
$H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy is thus used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes. This is the Min-Entropy (Eq. 5), i.e. the value in the limit $\alpha\to\infty$.
$H_{\infty}(X)\equiv\lim_{\alpha\to\infty}H_{\alpha}(X)\,=\,-\log\Bigl{(}\max_{x\in\mathcal{X}}p(x)\Bigr{)}.$ (5)
@@ 3.1 Threat Model and Assumptions
We consider an adversary whose primary objective is to compromise sensor-based key generation or authentication schemes by exploiting the inherent biases or predictable patterns within sensor data. We assume the adversary has full or near-full knowledge of the probability distribution of sensor outputs. They may obtain this through statistical modeling of publicly available datasets, prolonged observation of target users, or eavesdropping on sensor data streams.
We exclude capabilities such as fault-injection and other hardware attacks, e.g. see [34]. While such attacks could further reduce the effective entropy space—say, by inducing errors in the output values of sensing hardware—we regard this out-of-scope for this work. Furthermore, rather than brute-forcing all possible sensor outputs, we assume a more strategic approach. The adversary uses its knowledge of the sensor’s distribution to prioritise the most probable outcomes. By iterating guesses in order of descending probability, they minimise their expected effort. This threat model thus represents an adversary who can capitalise on statistical biases in sensor data without directly comprising the device physically. Our goal is to evaluate whether sensor data distributions—even aggregated from diverse users—offer sufficient entropy to resist knowledge-driven attacks.
@@ 3.2 High-level Methodology
    We merge sensor readings into a single, global distribution for each sensor modality in each dataset. For sensors that are inherently discrete or quantised (e.g., integer output ranges), we simply count occurrences. For sensors that produce (quasi-)continuous values, we rely on quantisation using Freedman–Diaconis binning to partition the output space and approximate an empirical probability mass function.
@@
  • University of Sussex–Huawei Locomotion (SHL) [40, 41]: sampled at 100 Hz from an Huawei Mate 9 smartphone. The publicly available SHL Preview dataset is used, comprising three recording-days per user (59 hours of data in total). To scope this study, we use the dataset from the handheld mobile phone as a good fit with related work (see §2).
@@ 4 Entropy Analysis
In this section, we analyse the intrinsic entropy of sensor data under the threat model described in §3. We begin by discussing the challenges in quantising naturally continuous sensor values for discrete-entropy calculations, then present our findings for our single- and multi-sensor analyses.
@@ 4.1 Pre-processing
A crucial, yet underexplored, issue in prior work (e.g. [22, 26, 27, 32]) is how to convert inherently continuous sensor outputs into suitable discrete values for entropy estimation. For example, Shannon and min-entropy, as defined in Eqs. 3 and 5, rely on discrete random variables. Physical quantities such as linear acceleration or angular velocity are continuous in nature, even though modern sensors employ internal analog-to-digital conversion with a finite resolution. Yet, a sensor’s advertised resolution (e.g. 12 bits for the widely used Bosch BMA mobile accelerometer [43]) does not imply uniform coverage across its range. Everyday usage introduces biases and clustering, resulting in some measurements occurring far more frequently than others. For instance, UCI-HAR data shows accelerometer readings concentrated in certain areas, and approximately 60% of gyroscope readings hover near zero (see Figure 2). Such skew and bias radically diminishes entropy compared to uniformly distributed values.
@@ 4.2 Single Sensors
Given the biases discussed above, it is inevitable that some sensor readings will exhibit relatively high predictability. To quantify this, we calculate individual-sensor entropies across multiple datasets. The results are given in Table 1. For multi-dimensional modalities (e.g. triaxial accelerometer or gyroscope), these are split into separate axes following Voris et al. [22]. We note that, in the Relay dataset, the data for individual $x$, $y$ and $z$ components are not given for the accelerometer, gyroscope, and magnetometer sensors. Rather, the authors have already preprocessed triaxial data into its vector magnitudes, i.e. $\textbf{v}=\sqrt{x^{2}+y^{2}+z^{2}}$. We give this as “X.Mag” for a given sensor X. For completeness, we compute the magnitude ourselves for other datasets, where applicable, and report the entropy values for this new synthetic modality.
@@
Table 1: Single-sensor entropy values (in bits) for each dataset. Grey cells denote unavailable data.